OpenStudy (anonymous):

what is the product of (z+5) cubed?

OpenStudy (anonymous):

Expand using the binomial theorem

ganeshie8 (ganeshie8):

did they teach u binomial theorem yet ?

ganeshie8 (ganeshie8):

if so, u have the power !

OpenStudy (sumi29):

Use the binomial theorem: In the general expression for the theorem, put n=3. Alternatively, for ease of use, try the Pascal's triangle to get the coefficients.

OpenStudy (anonymous):

Alternatively, for ease of use, try smashing your monitor against the wall several times

OpenStudy (sumi29):

Jack17: That's probably the best way....lol.

ganeshie8 (ganeshie8):

i get that feeling several times a day lol ;p

OpenStudy (unklerhaukus):

\[(z+5)^3=(z+5)^2\times(z+5)\\\qquad\qquad=(z^2+10z+25)\times(z+5)\\\qquad\qquad=\big((z^2+10z+25)\times z\big)+\big((z^2+10z+25)\times5\big)\]

OpenStudy (anonymous):

\[\left( a+b \right)^{3}=a ^{3}+b ^{3}+3ab \left( a+b \right)\] \[\left( z+5 \right)^{3}=z ^{3}+5 ^{3}+3*z*5\left( z+5 \right)\] you can solve it.